im having problems with this equation 7t/2-3=3t-2/3 i alot of different ways but for some reason when i check my answer i never get the same thing on both sides?
Answer by ptaylor(2198)

(7t/2)-3=3t-2/3 the LCM is 6; multiply both sides by 6
(7t/2)(6)-3(6)=3t(6)-6(2/3) clear fractions
21t-18=18t-4 subtract 18t and add 18 to both sides
21t-18t+18-18=18t-18t-4+18 collect like terms
3t=14
t=4.6666666---------------------------answer
ck by using calculator on original eq.

Hope this helps------ptaylor